A tax preparation company offers to advance you your $500 tax refund for a fee of $45. If you waited for the tax refund, you would get it within 90 days. So you are effectively borrowing $500 for 90 days and paying a fee (interest) of $45. Effectively, what simple annual interest rate are you paying? Assume a 365 day year. Enter your answer as a percent, rounded to the next whole number.
Answer #1

We know the formula for simple interest, that is:

I = P \times \frac{R}{100}\times T

Where,

I = Simple Interest paid

P = Principal amount

R = Rate of interest

T = Time in years.

Let us put the values in the formula.

45 = 500 \times \frac{R}{100}\times \frac{90}{365}

45 = \frac{R}{100}\times 123.2876712..

R = \frac{45}{1.2328767..}

R = 36.5\%

OR

R = 37\% ..............................Rounding it to the next whole number.